OpenStudy (anonymous):

cscx=2 from 0 to pi how do find values of x??

OpenStudy (anonymous):

Invert both sides and sin(x)=1/2?

OpenStudy (anonymous):

csc(x) = 1/sin(x) So: 1/sin(x) = y sin(x) = 1/y asin(sin(x)) = asin(1/y) = x

OpenStudy (anonymous):

There will be two solutions: one in quadrant one and one in quadrant two.

OpenStudy (anonymous):

Hmm, how do you find the second solution?

OpenStudy (anonymous):

-(x+pi), no?

OpenStudy (anonymous):

Unit circle helps. 

OpenStudy (anonymous):

x2=π-x1

OpenStudy (anonymous):

(1/sin(x))=2 then solve for sin(x)=1/2 x=arcsin(1/2)